HUBBLE SPACE TELESCOPE DAILY REPORT #814 PERIOD COVERED: 0800 EST Daily Status Report as of 033/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed PC 3230 (PC PSF and Focus Monitoring 2) The Planetary Camera was used to take two images of ninth magnitude standard star SAO43906 to monitor the point spread function and the focus changes caused by desorption within the WFPC system. Both image had to be filed as being bad because the point spread function of the target star fell on the Baum spot. 1.2 Completed One Set of WFC 4713 (WF/PC Calibration: WFC Internal Delta Flats) and Initiate a Second. The Wide Field Camera was used to obtain seven internal flat images. These observations will be used for the restoration of Earth flats, which were altered due to changes in quantum efficiency which occurred after decontamination operations. All of the images were received and displayed in the OSS. The first image was reported to be saturated in the corners, and was filed as being bad. 1.3 Completed GHRS/2 3961 (Elemental Abundance in Early Type Stars) The High Resolution Spectrograph was used to make high signal to noise ratio spectral observations of the chemically peculiar B class star Chi Lupie (AKA, HD141556). A total of twenty observations were received and displayed in the OSS. There was a carrousel reset during the reconfiguration for the deflection calibration. The following observations were wave length calibrations and observations of the target. They were reported to look okay, although part of one observation was idled and restarted. It showed broader absorption lines than the other groups. An HSTAR was written. 1.4 Completed WFC 3172 (WF/PC SV Observations: Internal Monitor Proposal) The Wide Field Camera was used to take three internal bias observations. No external targets were used. All three observations were received and displayed in the OSS. They all were reported to look okay. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 5 Successful : 5 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 9 Successful : 7 (FHST Failures 032/1833Z, and 033/1145Z) 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 4713 (WF/PC Calibration: WFC Internal Delta Flats) 3.2 Complete FOC 4107 (QSOs Suitable for Subsequent Observations of He II 304) 3.3 Complete PC 4593 (Positioning of Geminga Optial Counterpart) 3.4 Complete WFC/PC 3172 (WF/PC SV Observations: Internal Monitor Proposal) 3.5 Complete FOC/96 3801 (QSOs Suitable for Subsequent Observations) 3.6 Complete FOS/BL 2424 (Quasar Absorption Line Survey: Observations FOS) 3.7 Complete WFC 3917 (WFC:HST Medium Deep Survey Part 1) 3.8 Complete One Set of PC 4714 (WF/PC Calibration:Internal Delta Flats) and Initiate a Second. 3.9 Complete FOS/BL 3550 (Absorption Lines from Gas in Cooling Flow) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END- ----- End Included Message -----